OpenStudy (samsan9):

identify the conic section from its equation. (x-5)^2/4+(y+3)^2/16=1

OpenStudy (imstuck):

this is an ellipse

OpenStudy (samsan9):

are all the ones that look like this equation ellipses?

OpenStudy (samsan9):

@IMStuck

OpenStudy (imstuck):

Ok, ellipses are x^2 + y^2 over a and b=1, and the a is always bigger than the b. That tells you which vertices are the major and which are the minor. Hyperbolas are x^2 - y^2=1. Parabolas are either an x^2 or a y^2, but not both in the same equation. Do you have more of these?

OpenStudy (imstuck):

If they look like that with an x^2 first and it is added to a Y^2, then yes, it is an ellipse.

OpenStudy (samsan9):

yes i do um there is one that says X^2+y^2+6x-16y+48=0

OpenStudy (imstuck):

Elliipse.

OpenStudy (imstuck):

Nononono circle I think. Hold on just one sec!!!!

OpenStudy (samsan9):

ok

OpenStudy (imstuck):

yes it is a circle. I had to complete the square and get it into standard form which is:\[(x+3)^{2}+(y-8)^{2}=25\]That is a circle with center (-3,8) and a radius of 5.

OpenStudy (imstuck):

If it was an ellipse it would HAVE to equal 1 and this one does not work out that way.

OpenStudy (samsan9):

what about one that says y^2+6y-2x+13=0

OpenStudy (imstuck):

Here are a couple of rules to keep in the back of your mind regarding quadratic equations: In the second degree equation\[Ax ^{2}+By ^{2}+Cx+Dy+E=0\]j(which is the format of your equation exactly!): 1. If A=B, the equation may define a circle. 2. If A and B have the same sign and A does NOT equal B, the equation may define an ellipse.

OpenStudy (imstuck):

Here your A=B so it cannot be an ellipse.

OpenStudy (imstuck):

The one you just asked about has a y^2 only. That makes it a parabola in the form \[y ^{2}=4px\]

OpenStudy (imstuck):

It opens to the right and the focus lies on the x axis and the directrix is an x= ??? line.

OpenStudy (imstuck):

We could solve those if you need to eventually...

OpenStudy (imstuck):

I love these things. They were SO hard for me to learn; I literally spent HOURS mastering them.

OpenStudy (imstuck):

Anyway, enough of that! ; ) Tag me if you need me again!

OpenStudy (samsan9):

ok thank you :D